# Benchmarking Large Vision-Language Models on Fine-Grained Image Tasks: A Comprehensive Evaluation

### v1

@@ Abstract
Recent advancements in Large Vision-Language Models (LVLMs) have demonstrated remarkable multimodal perception capabilities, garnering significant attention. While numerous evaluation studies have emerged, assessing LVLMs both holistically and on specialized tasks, fine-grained image tasks—fundamental to computer vision—remain largely unexplored. To fill this gap, we introduce a comprehensive fine-grained evaluation benchmark, i.e.FG-BMK, comprising 3.49 million questions and 3.32 million images. Our evaluation systematically examines LVLMs from both human-oriented and machine-oriented perspectives, focusing on their semantic recognition and fine-grained feature representation capabilities. Through extensive experiments on eight representative LVLMs/VLMs, we uncover key findings regarding the influence of training paradigms, modality alignment, perturbation susceptibility, and fine-grained category reasoning on task performance. This work provides critical insights into the limitations of current LVLMs and offers guidance for future data construction and model design in the development of more advanced LVLMs. Our code is open-source and available at https://github.com/SEU-VIPGroup/FG-BMK.
@@ 1 Introduction
To address this gap, we propose a comprehensive fine-grained evaluation benchmark, termed FG-BMK. This benchmark includes 3.49 million questions and 3.32 million images, providing a robust test bed for evaluating LVLMs. FG-BMK incorporates two evaluation paradigms: human-oriented and machine-oriented. The human-oriented evaluation employs dialogue-like interactions to assess a model’s ability to understand and respond to fine-grained visual queries in a conversational context. The machine-oriented evaluation focuses on two core fine-grained vision tasks—image retrieval and recognition—to directly measure the feature representation capabilities of LVLMs. Together, these evaluations enable a comprehensive assessment of LVLMs’ fine-grained feature representation and semantic recognition abilities.
@@ 2 Related Work
Large Language Models (LLMs) like GPT-4 [1] have achieved significant advancements in text comprehension, reasoning, and generation. Building on this foundation, Large Vision-Language Models (LVLMs) have emerged, demonstrating remarkable reasoning and perception capabilities across various tasks. Efforts to enhance LVLM performance have taken diverse approaches. BLIP [20] leverages noisy web data with bootstrapped captions, while LLaVA [22] utilizes GPT-4-generated instruction-following data to improve visual-linguistic understanding. BLIP-2 [19] focuses on integrating frozen image and text encoders, and Qwen-VL [2] employs a three-stage training pipeline with a generative loss. InternVL [7] scales up the vision encoder to 6 billion parameters to address parameter size disparities. Despite their success across various tasks, these LVLMs’ capabilities in fine-grained visual tasks remain underexplored. Our work seeks to fill this gap by evaluating their performance in this challenging domain, providing deeper insights into their potential and limitations.
@@ 4.1 Models under Evaluation
Given the diverse landscape of existing LVLMs, we select seven widely-used representative LVLMs and one purely visual model with varying training strategies, as shown in Table 1. To enable clear comparisons and isolate the factors influencing performance, we use earlier versions of models from each family, where their distinct characteristics are more evident and less affected by additional tricks or complex modifications. Further details about the evaluated models can be found in Appendix B.
@@ 4.2 Human-oriented Evaluation
Table 1: Training strategies of the evaluated models. “DINOv2" is a purely visual model. “Con" denotes contrastive loss, “Gen" generative loss, “Mat" image-text matching loss, “Rec" reconstruction loss used in BEiT3, and “Dis" distillation loss used in DINOv2.
@@
Figure 2: Results of LLaVa [22] on true/false and multiple-choice questions across different levels of granularity. The $x$-axis denotes the granularity of the recognition questions.
As shown in Figure 2, taking LLaVA [22] as an example, its accuracy in answering true/false and multiple-choice questions declines as the granularity becomes finer. At the class level (e.g., “Is the class of the object in this image an Insecta/Aves?”), the model achieves 99.97% accuracy on multiple-choice questions and 98.10% on true/false questions. [^1] When questions are relatively simple, LVLMs achieve very high accuracy. The higher accuracy on multiple-choice questions compared to true/false questions may result from randomness. However, as the granularity narrows to the genus level, where negative categories are drawn from different genera within the same class (e.g., “Is the object in this image an albatross or a gull?”), the model’s accuracy on multiple-choice questions drops to 66.78%, a decrease of 33.19%. When the granularity is further refined to the species level, where negative categories are from different species within the same genus (e.g., “Is the object in this image a black-footed albatross/Laysan albatross?”), the accuracy drops further to 59.04% on true/false questions and 42.58% on multiple-choice questions. This demonstrates that the model struggles to distinguish between closely related species. The results for other LVLMs exhibit similar trends to those observed for LLaVA. Additional examples of multiple-choice and true/false questions can be found in Appendix C.1.
@@
To examine whether LVLMs exhibit knowledge bias in recognizing different fine-grained categories, we ranked their accuracy in answering true/false questions for each fine-grained category. As shown in Figure 3, using LLaVA [22] as an example, the model demonstrates highly inconsistent recognition abilities across categories, achieving nearly 90% accuracy for some while dropping to approximately 20% for others.
@@
belly color 54.79 back color 41.90 bill color 41.44 breast color 49.56
crown color 48.71 eye color 69.27 forehead color 47.03 leg color 35.37
@@
tail pattern 23.19 wing pattern 29.67
@@
bill shape 1.39 shape 18.59 tail shape 9.89 wing shape 26.34
@@
bill length 15.71 size 49.47
@@
As shown in Table 2 of the paper and Table 8 in the appendix, LVLMs exhibit relatively stronger recognition abilities for attributes like color and size compared to others. For example, the average accuracy for color recognition is 44.43% for LLaVA [22] and 38.14% for BLIP2 [19], whereas for shape recognition, the accuracy drops significantly to 14.05% and 25.05%, respectively. Notably, no attribute achieved an accuracy above 70%, and several attributes scored as low as 10%, highlighting significant room for improvement in LVLMs’ fine-grained attribute recognition.
We also observe model-specific differences in attribute recognition. For instance, while LLaVA struggles more with color recognition compared to size, BLIP2 shows the opposite trend. Additionally, although LLaVA surpasses BLIP2 in color recognition, it underperforms in recognizing attributes like length. Detailed results of the attribute recognition task are provided in Appendix C.3.
@@ 4.3 Machine-oriented Evaluation
Regarding the amount of training data, as shown in Figure 7, the vision encoder of EVA-CLIP, trained on over 2 billion data samples, does not outperform DINOv2, which was trained on 142 million samples, in fine-grained classification and retrieval tasks. We attribute this difference to the quality of the training data, that DINOv2’s dataset is carefully curated from a large pool of data, whereas EVA-CLIP relies on crawled web data. A similar trend is observed when comparing DINOv2 with InternVL, which was trained on over 6 billion web data samples. This indicates that simply increasing the scale of training data, without considering its quality, offers limited improvement in fine-grained discriminability of visual features.
@@
As shown in Table 5, perturbations to the visual features of EVA-CLIP result in a significant drop in classification accuracy on the fine-grained dataset CUB-200-2011 (from 88.95% to 24.94%). In contrast, on the generic dataset CIFAR-100 [18], the accuracy declines more modestly (from 93.05% to 50.76%). Similar trends are observed for CoCa and DINOv2, indicating that fine-grained tasks are more susceptible to perturbations than generic tasks. We attribute this vulnerability to the coarse-grained and noisy data used during training, which limits the discriminability of features across fine-grained categories, making it challenging to distinguish between them when visual features are perturbed.
@@ 5 Concluding Remarks
Despite FG-BMK covering most fine-grained image tasks, it is still impossible to cover every fine-grained scenario in real-world applications. For instance, some specific domains, such as fine-grained medical imaging tasks or fine-grained remote sensing change detection tasks, might not be fully represented. While, we believe that FG-BMK provides the most practical setting for evaluating LVLMs in most use cases, and the conclusions drawn from this benchmark should be applicable when generalizing to real-world fine-grained visual tasks.
@@ A.1 Evaluation Task Details
In Section 3.1, we have described each evaluation task. Here, we provide further details. In the Knowledge Bias Estimation task, to uncover potential knowledge biases across different fine-grained categories, we pair each fine-grained category label with images from other subcategories within the same super-category to generate negative samples for true/false questions. For each fine-grained category label, we calculate the LVLM’s accuracy on all coresponding true/false questions as a measure of its understanding of that category’s knowledge.
@@ Appendix B Evaluated Models
As shown in Table 7, we evaluate seven representative LVLMs and one purely visual model, each of which employs a distinctive training recipes, including variations in vision encoder, language model, training losses and data.
Table 7: Configurations of the evaluated models. “DINOv2-L” is a purely visual model. “–” indicates the model is trained from scratch. “Con” stands for the contrastive loss, “Gen” for the generative loss, “Mat” for the image-text matching loss, “Rec” for the reconstruction loss as used in BEiT3 [40], and “Dis” for the distillation loss as applied in DINOv2 [31].
Model Component Loss Function Training Data
Vision Model Language Model Con Gen Mat Rec Dis $<$ 100M 100M $\sim$ 1B $>$ 1B
@@ C.1 Results of Hierarchical Granularity Recognition
Figure 2 shows LLaVA’s [22] accuracy in answering true/false and multiple-choice questions within hierarchical granularity recognition task on CUB-200-2011 dataset. In Figure 11, Figure 11 and Figure 11, we present additional results for LLaVA and InternVL [7] on CUB-200-2011 [39] and iNat [38] datasets. As shown in the experiments, the accuracy of all models decreases as the granularity becomes finer. When the granularity level reaches the finest level, the models struggle to distinguish between closely related species.
(a) LLaVA on iNat2021
(b) InternVL on iNat2021
(c) InternVL on CUB-200-2011
Figure 11: Results of LLaVA [22] and InternVL [7] on true/false and multiple-choice questions across different levels of granularity on CUB-200-2011 [39] and iNat2021 [38] dataset. The x-axis denotes the granularity of the recognition questions.
@@ C.2 Results of Knowledge Bias Estimation
In Figure 3, we observe that LVLMs exhibit highly inconsistent recognition abilities across categories. We also conduct experiments with LLaVA, InternVL, and BLIP2 on fine-grained datasets such as Stanford Dog, Flowers102, and Aircraft. As shown in Figure 12, all LVLMs display similar trends, indicating inconsistent recognition abilities across categories. However, after fine-tuned on datasets with balanced occurrences of fine-grained categories, all LVLMs demonstrate remarkable recognition abilities across all fine-grained categories.
(a) LLaVA on Flowers102
(b) InternVL on Stanford Dog
(c) BLIP2 on Aircraft
Figure 12: Comparison of the original and fine-tuned LVLMs on occurrence-balanced fine-grained categories. True/false question accuracy for each category is ranked, with blue dots representing the original model and yellow dots the fine-tuned model.
@@ C.3 Results of Attribute Recognition
Table 2 and Table 8 shows the attribute recognition accuracy of LLaVA and BLIP2 on the CUB-200-2011 dataset. The results of InternVL and Qwen are shown in Table

### v2

@@ Abstract
Recent advancements in Large Vision-Language Models (LVLMs) have demonstrated remarkable multimodal perception capabilities, garnering significant attention. While numerous evaluation studies have emerged, assessing LVLMs both holistically and on specialized tasks, fine-grained image tasks—fundamental to computer vision—remain largely unexplored. To fill this gap, we introduce a comprehensive fine-grained evaluation benchmark, i.e.FG-BMK, comprising 1.01 million questions and 0.33 million images. Our evaluation systematically examines LVLMs from both human-oriented and machine-oriented perspectives, focusing on their semantic recognition and fine-grained feature representation capabilities. Through extensive experiments on twelve representative LVLMs/VLMs, we uncover key findings regarding the influence of training paradigms, modality alignment, perturbation susceptibility, and fine-grained category reasoning on task performance. This work provides critical insights into the limitations of current LVLMs and offers guidance for future data construction and model design in the development of more advanced LVLMs. Our code is open-source and available at https://anonymous.4open.science/r/FG-BMK-7B51.
@@ 1 Introduction
To address this gap, we propose a comprehensive fine-grained evaluation benchmark, termed FG-BMK. This benchmark includes 1.01 million questions and 0.33 million images, providing a robust test bed for evaluating LVLMs. FG-BMK incorporates two evaluation paradigms: human-oriented and machine-oriented. The human-oriented evaluation employs dialogue-like interactions to assess a model’s ability to understand and respond to fine-grained visual queries in a conversational context. The machine-oriented evaluation focuses on two core fine-grained vision tasks—image retrieval and recognition—to directly measure the feature representation capabilities of LVLMs. Together, these evaluations enable a comprehensive assessment of LVLMs’ fine-grained feature representation and semantic recognition abilities.
@@ 2 Related Work
Large Language Models (LLMs) like GPT-4 [1] have achieved significant advancements in text comprehension, reasoning, and generation. Building on this foundation, Large Vision-Language Models (LVLMs) have emerged, demonstrating remarkable reasoning and perception capabilities across various tasks. Efforts to enhance LVLM performance have taken diverse approaches. BLIP [22] leverages noisy web data with bootstrapped captions, while LLaVA [24] utilizes GPT-4-generated instruction-following data to improve visual-linguistic understanding. BLIP-2 [21] focuses on integrating frozen image and text encoders, and Qwen2.5-VL [2] employs dynamic-resolution strategy to preserve native image resolution. InternVL3 [55] adopts a unified pre-training approach over multimodal data to improve effectiveness. Despite their success across various tasks, these LVLMs’ capabilities in fine-grained visual tasks remain underexplored. Our work seeks to fill this gap by evaluating their performance in this challenging domain, providing deeper insights into their potential and limitations.
@@ 4.1 Models under Evaluation
Given the diverse landscape of existing LVLMs, we select nine widely-used open-source LVLMs, two closed-source models (GPT-4o-1120 [1] and Gemini-2.0-flash [13]) and one purely visual model with varying training strategies, as shown in Table 1. To enable clear comparisons and isolate the factors influencing performance, we use earlier versions of models from each family in machine-oriented evaluation, where their distinct characteristics are more evident and less affected by additional tricks or complex modifications. Further details about the evaluated models can be found in Appendix B.
@@ 4.2 Human-oriented Evaluation
Table 1: Training strategies of the open-source evaluated models. “DINOv2" is a purely visual model. “Con" denotes contrastive loss, “Gen" generative loss, “Mat" image-text matching loss, “Rec" reconstruction loss used in BEiT3, and “Dis" distillation loss used in DINOv2.
@@
Figure 2: Results of InternVL3 [55] on true/false and multiple-choice questions across different levels of granularity on the CUB-200-2011 [41] dataset. The $x$-axis denotes the granularity of the recognition questions.
As shown in Figure 2, taking InternVL3 [55] as an example, its accuracy in answering true/false and multiple-choice questions declines as the granularity becomes finer. At the class level (e.g., “Is the class of the object in this image an Insecta/Aves?”), the model achieves 99.76% accuracy on multiple-choice questions and 99.77% on true/false questions. [^1] When questions are relatively simple, LVLMs achieve very high accuracy. The higher accuracy on multiple-choice questions compared to true/false questions may result from randomness. However, as the granularity narrows to the genus level, where negative categories are drawn from different genera within the same class (e.g., “Is the object in this image an albatross or a gull?”), the model’s accuracy on multiple-choice questions drops to 90.75%, a decrease of 9.01%. When the granularity is further refined to the species level, where negative categories are from different species within the same genus (e.g., “Is the object in this image a black-footed albatross/Laysan albatross?”), the accuracy drops further to 62.48% on true/false questions and 61.18% on multiple-choice questions. This demonstrates that the model struggles to distinguish between closely related species. The results for other LVLMs exhibit similar trends to those observed for InternVL3. Additional examples of multiple-choice and true/false questions can be found in Appendix C.1.
@@
To examine whether LVLMs exhibit knowledge bias in recognizing different fine-grained categories, we ranked their accuracy in answering true/false questions for each fine-grained category. As shown in Figure 3, using LLaVA [24] as an example, the model demonstrates highly inconsistent recognition abilities across categories, achieving nearly 90% accuracy for some while dropping to approximately 30% for others.
@@
Table 2: Attribute recognition accuracy of InternVL3 [55] on the CUB-200-2011 [41] dataset (values in parentheses represent the average accuracy for each attribute).
@@
belly color 58.49 back color 34.98 bill color 51.31 breast color 54.25
crown color 55.30 eye color 84.59 forehead color 53.32 leg color 44.01
nape color 39.24 throat color 52.77 under tail color 34.69 underparts color 56.20
upper tail color 37.30 upperparts color 28.75 wing color 30.16 primary color 43.05
@@
back pattern 40.94 belly pattern 68.13 breast pattern 65.12 head pattern 35.92
tail pattern 41.64 wing pattern 49.04
@@
bill shape 37.61 shape 52.37 tail shape 10.42 wing shape 23.39
@@
bill length 71.03 size 52.55
@@
As shown in Table 2 of the paper and Table 12 in the appendix, LVLMs exhibit relatively stronger recognition abilities for attributes like pattern, size and length compared to others. For example, InternVL3 [55] and Qwen2.5-VL [3] achieve 50.13% and 45.12% average accuracy for pattern recognition, but only 30.95% and 29.30% for shape recognition. Notably, only a few attributes achieved an accuracy above 70%, and several attributes scored as low as 10%, highlighting significant room for improvement in LVLMs’ fine-grained attribute recognition.
We also observe model-specific differences in attribute recognition. For instance, while InternVL3 struggles more with pattern recognition compared to size, Gemini-2.0-flash [13] shows the opposite trend. Additionally, we find that the latest LVLMs show significant improvements in pattern and length recognition, while their advancements in color and shape recognition are more limited. Detailed results of the attribute recognition task are provided in Appendix C.3.
@@ 4.3 Machine-oriented Evaluation
Regarding the amount of training data, as shown in Figure 7, the vision encoder of EVA-CLIP, trained on over 2 billion data samples, does not outperform DINOv2, which was trained on 142 million samples, in fine-grained classification and retrieval tasks. We attribute this difference to the quality of the training data, that DINOv2’s dataset is carefully curated from a large pool of data, whereas EVA-CLIP relies on crawled web data. A similar trend is observed when comparing DINOv2 with InternVL (6B samples), suggesting that simply increasing the scale of training data, without considering its quality, offers limited improvement in fine-grained discriminability of visual features.
@@
As shown in Table 5, perturbing the visual features of EVA-CLIP causes a significant drop in classification accuracy on the fine-grained dataset CUB-200-2011 (from 88.95% to 24.94%). In contrast, on the generic dataset CIFAR-100 [20], the accuracy declines more modestly (from 93.05% to 50.76%). Similar trends are observed for CoCa and DINOv2, indicating that fine-grained tasks are more susceptible to perturbations than generic tasks. We attribute this vulnerability to the coarse-grained and noisy data used during training, which limits the discriminability of features across fine-grained categories, making it challenging to distinguish between them when visual features are perturbed.
@@ 5 Concluding Remarks
Despite FG-BMK covers most fine-grained image tasks, it is still impossible to cover every fine-grained scenario in real-world applications, such as medical imaging or remote sensing. Nevertheless, we believe that FG-BMK provides the most practical setting for evaluating LVLMs in most use cases, and the conclusions drawn from this benchmark should be applicable to real-world fine-grained visual tasks.
@@ A.1 Evaluation Task Details
In Section 3.1, we have described each evaluation task. Here, we provide further details. In the Knowledge Bias Estimation task, to uncover potential knowledge biases across different fine-grained categories, we pair each image with its corresponding fine-grained label to generate positive samples for true/false questions. For constructing negative samples, each image is paired with a single fine-grained label randomly selected from other subcategories within the same super-category. For each fine-grained category, we calculate the LVLM’s accuracy on all coresponding true/false questions as a measure of its understanding of that category’s knowledge.
@@ Appendix B Evaluated Models
As shown in Table 7, we select nine widely-used open-source LVLMs, two closed-source models (GPT-4o [1] and Gemini-2.0-flash [13]) and one purely visual model, each of which employs a distinctive training recipes, including variations in vision encoder, language model, training losses and data.
Table 7: Configurations of the evaluated models. “DINOv2-L” is a purely visual model. “Con” stands for the contrastive loss, “Gen” for the generative loss, “Mat” for the image-text matching loss, “Rec” for the reconstruction loss as used in BEiT3 [42], and “Dis” for the distillation loss as applied in DINOv2 [33].
Model Component Loss Function
Vision Model Language Model Con Gen Mat Rec Dis
InternV3-7B InternViT-L Qwen2.5-7B
@@
  • InternVL3 [55] adopts a unified pre-training approach over both multimodal and pure-text data, enhanced by variable visual position encoding (V2PE) and advanced post-training strategies for improved scalability and effectiveness.
@@
  • Qwen2.5-VL [3] combines dynamic-resolution Vision Transformer with Window Attention to reduce computational cost while preserving native image resolution.
@@
For all our evaluated model, we follow their official configurations to run the inference. We set the temperature of all open-source models to 0, while keeping the default for closed-source APIs.
@@ C.1 Results of Hierarchical Granularity Recognition
Figure 2 shows InternVL3’s [55] accuracy in answering true/false and multiple-choice questions within hierarchical granularity recognition task on CUB-200-2011 dataset. In Figure 11, we present additional results for GPT-4o [1], Gemini-2.0-flash [13], Qwen2.5-VL [3], LLaVA [24] and InternVL [8] on CUB-200-2011 [41] and iNat2021 [40] datasets. As shown in the experiments, the accuracy of all models decreases as the granularity becomes finer. When the granularity level reaches the finest level, the models struggle to distinguish between closely related species.
(a) GPT-4o on iNat2021
(b) Genimi-2.0-flash on iNat2021
(c) LLaVA on iNat2021
(d) InternVL on iNat2021
(e) LLaVA on CUB-200-2011
(f) Qwen2.5-VL on CUB-200-2011
(g) InternVL on CUB-200-2011
Figure 11: Results of GPT-4o [1], Gemini-2.0-flash [13], Qwen2.5-VL [3], LLaVA [24] and InternVL [8] on true/false and multiple-choice questions across different levels of granularity on CUB-200-2011 [41] and iNat2021 [40] dataset. The x-axis denotes the granularity of the recognition questions.
@@ C.2 Results of Knowledge Bias Estimation
In Figure 3, we observe that LLaVA exhibit highly inconsistent recognition abilities across categories. We also conduct experiments with Qwen2.5-VL, GPT-4o, and Gemini-2.0-flash on fine-grained datasets such as Aircraft [29], Flowers102 [32] and Stanford Dogs [18]. As shown in Figure 12 and Figure 13, all LVLMs display similar trends, indicating inconsistent recognition abilities across fine-grained categories. However, after fine-tuned on datasets with balanced occurrences of fine-grained categories, LVLMs demonstrate remarkable recognition abilities across all fine-grained categories.
To construct datasets with balanced occurrences of fine-grained categories, we select an equal number of images from each category. Then we generate the same number of true/false questions for each fine-grained category, thereby fine-tuning the LVLMs in a way that each category receives balanced representation.
(a) GPT-4o on Flowers102
(b) Gemini-2.0-flash on Stanford Dog
Figure 12: Knowledge bias estimation results of two closed-source models. True/false question accuracy for each category is ranked, with blue dots representing the original model.
Figure 13: Comparison of the original and fine-tuned Qwen2.5-VL [3] models on occurrence-balanced fine-grained aircraft categories. True/false question accuracy for each category is ranked, with blue dots representing the original model and yellow dots the fine-tuned model.
@@ C.3 Results of Attribute Recognition
Table 2 and Table 12 shows the attribute recognition accuracy of InternVL3 and Qwen2.5-VL on the CUB-200-2011 dataset. The results of LLaVA, BLIP2, InternVL and Gemini-2.0-flash are shown in Table 8, Table 9, Table 10, and Table 11.
Color Attribute (44.34)
belly color 54.79 back color 41.90 bill color 41.44 breast color 49.56
crown color 48.71 eye color 69.27 forehead color 47.03 leg color 35.37
Pattern Attribute (23.69)
tail pattern 23.19 wing pattern 29.67
Shape Attribute (14.05)
bill shape 1.39 shape 18.59 tail shape 9.89 wing shape 26.34
Length Attribute (15.71) Size Attribute (49.47)
bill length 15.71 size 49.47
@@
Table 11: Attribute recognition accuracy of Gemini-2.0-flash [13] on the CUB-200-2011 [41] dataset (values in parentheses represent the average accuracy for each attribute).
Color Attribute (47.22)
belly color 62.09 back color 36.51 bill color 52.31 breast color 56.01
crown color 56.44 eye color 59.57 forehead color 53.55 leg color 40.66
nape color 40.40 throat color 60.23 under tail color 40.60 underparts color 59.65
upper tail color 39.99 upperparts color 29.66 wing color 29.21 primary color 38.69
Pattern Attribute (56.14)
back pattern 56.26 belly pattern 70.51 breast pattern 66.89 head pattern 39.56
tail pattern 52.33 wing pattern 51.26
Shape Attribute (48.75)
bill shape 61.62 shape 68.20 tail shape 32.13 wing shape 33.04
Length Attribute (71.82) Size Attribute (52.72)
bill length 71.82 size 52.72
Table 12: Attribute recognition accuracy of Qwen2.5-VL [3] on the CUB-200-2011 [41] dataset (values in parentheses represent the average accuracy for each attribute).
Color Attribute (40.39)
belly color 51.11 back color 32.89 bill color 46.50 breast color 44.84
crown color 46.54 eye color 54.85 forehead color 44.57 leg color 37.79
nape color 36.49 throat color 40.74 under tail color 34.60 underparts color 50.20
upper tail color 34.92 upperparts color 27.20 wing color 26.03 primary color 36.96
Pattern Attribute (45.12)
back pattern 42.66 belly pattern 64.58 breast pattern 59.79 head pattern 14.57
tail pattern 45.04 wing pattern 44.11
Shape Attribute (29.30)
bill shape 15.30 shape 58.17 tail shape 5.63 wing shape 38.10
Length Attribute (63.20) Size Attribute (52.56)
bill length 63.20 size 52.56